a rectangular room is 25 feet narrower than its length. if 1250 square feet of carpet is needed to cover the entire room, what is the length of the room?
Answer by JulietG(1812)

Area = Length * Width
1250 = L * W
W = L - 25 [a rectangular room is 25 feet narrower than its length.]
Substitute the known value of W from the second equation into the first equation.
1250 = L * W
1250 = L * (L-25)
Divide each side by 25
L = 50
Therefore W = 25